1. The airbag is a safety device used in cars to cushion the passenger during a crash. it involves the following chemical reaction which is triggered by an impact.

2 Na N3(s) à ƒ ƒ ¢ † ’ 2 Na(s) + 3 N2(g)

Calculate the volume of N2 gas (measured at STP) that can be produced from 130.0 g of Na N3 (molar mass =65).

2. The compound lithium carbonate, used to treat manic depression, has the formula of Li2CO3. Calculate the number of formula units of lithium carbonate present in a 0.500g sample of lithium carbonate.
1
Expert's answer
2021-10-14T04:04:51-0400

Number of moles NaN3= 130/65= 2

Mole ratio NaN3:N2= 2:3

= 2×3/2= 3moles

At STP, 1 mole = 22.4L

3 × 22.4L = 67.2L